4x^2+16x+16 solve for x
Answer by drumwrrv(7)

4x^(2)+16x+16=0
Factor out the GCF of 4 from each term in the polynomial.
4(x^(2))+4(4x)+4(4)=0
Factor out the GCF of 4 from 4x^(2)+16x+16.
4(x^(2)+4x+4)=0
For a polynomial of the form x^(2)+bx+c, find two factors of c (4) that add up to b (4). In this problem 2*2=4 and 2+2=4, so insert 2 as the right hand term of one factor and 2 as the right-hand term of the other factor.
4(x+2)(x+2)=0
Combine the two common factors of (x+2) by adding the exponents.
4(x+2)^(2)=0
Divide both sides of the equation by 4. Dividing 0 by any non-zero number is 0.
(x+2)^(2)=0
Set each of the factors of the left-hand side of the equation equal to 0.
x+2=0
Since 2 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 2 from both sides.
x=-2